# 3D-GP-LMVIC: Learning-based Multi-View Image Coding with 3D Gaussian Geometric Priors

### v1

@@ Abstract
Multi-view image compression is vital for 3D-related applications. To effectively model correlations between views, existing methods typically predict disparity between two views on a 2D plane, which works well for small disparities, such as in stereo images, but struggles with larger disparities caused by significant view changes. To address this, we propose a novel approach: learning-based multi-view image coding with 3D Gaussian geometric priors (3D-GP-LMVIC). Our method leverages 3D Gaussian Splatting to derive geometric priors of the 3D scene, enabling more accurate disparity estimation across views within the compression model. Additionally, we introduce a depth map compression model to reduce redundancy in geometric information between views. A multi-view sequence ordering method is also proposed to enhance correlations between adjacent views. Experimental results demonstrate that 3D-GP-LMVIC surpasses both traditional and learning-based methods in performance, while maintaining fast encoding and decoding speed. The code is available at https://github.com/YujunHuang063/3D-GP-LMVIC.
@@ 1 Introductioin
The rapid advancement of 3D applications has led to an explosion of multi-view image data across various fields such as virtual reality (VR) (Anthes et al., 2016), augmented reality (AR) (Schmalstieg & Hollerer, 2016), 3D scene understanding (Dai et al., 2017), autonomous driving (Chen et al., 2017), and medical imaging (Hosseinian & Arefi, 2015). Applications like VR and AR, which rely on high-quality multi-view visual content to create immersive experiences, generate a massive volume of data that poses significant challenges for storage and transmission. This makes the development of efficient compression techniques crucial for managing the increasing data demands in these fields.
@@
Current standards, such as H.264-based MVC (Vetro et al., 2011) and H.265-based MV-HEVC (Hannuksela et al., 2015), have been developed to compress multi-view media by extending their respective base standards and exploiting redundancies across multiple views. These standards incorporate inter-view prediction, where blocks in one view are predicted based on corresponding blocks in neighboring view. Disparity estimation is employed to calculate positional differences of objects between views, aiding in the prediction of pixel values. However, these methods rely on manually designed modules, limiting the system’s ability to fully leverage end-to-end optimization. Consequently, the spatial correlation among views may not be fully exploited, leading to suboptimal compression performance.
@@
Learning-based single image compression has seen remarkable advancements (Ballé et al., 2017; 2018; Minnen et al., 2018), inspiring extensions of these methods to multi-view image coding (Deng et al., 2021; Lei et al., 2022; Zhang et al., 2023; Liu et al., 2024). A central challenge in these extensions lies in the accurate estimation of disparities across different views. For example, Deng et al. (2021; 2023) employ a simple 3x3 homography matrix for disparity estimation, which, while efficient, struggles with complex scene disparities. Alternatively, Ayzik & Avidan (2020); Huang et al. (2023) utilize patch matching method to align the reference view with the target view. This approach is effective for horizontal or vertical view shifts but falls short when addressing non-rigid deformations caused by view rotations. Similarly, Zhai et al. (2022) assume that disparity occurs only along the horizontal axis in their stereo matching method, which suffices for stereo images but is inadequate for more complex view transformations where disparity is not limited to the horizontal axis. Some methods leverage cross-attention mechanisms for implicit alignment (Wödlinger et al., 2022; Zhang et al., 2023; Liu et al., 2024). For instance, Zhang et al. (2023) enhance the target view’s representation by multiplying its query with the reference view’s key and value, effectively incorporating reference view features into the target view. However, these methods primarily establish correlations between two views on a 2D plane, without considering the 3D spatial relationships between the views and the objects being captured.
In addition, we are impressed by 3D Gaussian Splatting (3D-GS) (Kerbl et al., 2023), a novel technique for generating 3D representations from multi-view images and synthesizing novel views. This method achieves remarkable results by representing scenes as mixtures of small, colored Gaussians, capturing intricate geometric details and continuous depth and texture variations. The probabilistic nature of Gaussians allows for smooth, accurate modeling of complex surfaces, enabling the creation of realistic and detailed 3D environments. Additionally, the fast differentiable renderer inherent in 3D Gaussian Splatting ensures efficient real-time inference.
@@
Building on prior research, we propose a learning-based multi-view image coding framework incorporating 3D Gaussian geometric priors (3D-GP-LMVIC). This framework employs 3D-GS as a geometric prior to guide disparity estimation across views. Specifically, 3D-GS generates a depth map for each view, offering precise spatial location information for each pixel of the captured object. This spatial data facilitates accurate pixel-level correspondence across views, enabling the compression model to more effectively fuse features from reference views. Additionally, due to positional and angular disparities, images from different views may not fully overlap, and merging non-overlapping regions can introduce noise. To mitigate this, a mask derived from the 3D Gaussian geometric prior is designed to identify overlapping regions between views. Moreover, as depth maps are necessary during decoding, we introduce a depth map compression model to efficiently reduce geometric redundancy across views. This model incorporates a cross-view depth prediction module to capture geometric correlations. Recognizing the critical role of view correlation in redundancy reduction, we also propose a multi-view sequence ordering method to address the issue of low overlap between adjacent views in unordered sequences. This method defines a distance measure between pairs of views to guide the ordering of view sequences.
@@
  • We propose a learning-based multi-view image coding framework with 3D Gaussian geometric priors (3D-GP-LMVIC), which utilizes 3D Gaussian geometric priors for precise disparity estimation across views, thereby enhancing multi-view image compression efficiency. Additionally, we design a mask based on these priors to identify overlapping regions between views, effectively guiding the model to retain useful cross-view information.
@@
  • We also present a depth map compression model aimed at reducing geometric redundancy across views. Furthermore, a multi-view sequence ordering method is proposed to improve the correlation between adjacent views.
@@
  • Experimental results show that our framework surpasses both traditional and learning-based multi-view image coding methods in compression efficiency, while also providing fast encoding and decoding speed. Moreover, our disparity estimation method demonstrates greater visual accuracy compared to existing two-view disparity estimation methods.
@@ 2 Related Works
In recent years, learning-based image compression methods have integrated autoencoders with differentiable entropy models to enable end-to-end optimization of rate-distortion loss. Early works, such as Ballé et al. (2017; 2018), introduced generalized divisive normalization (GDN) (Ballé et al., 2016) and proposed factorized and hyperprior entropy models. Subsequent research (Minnen et al., 2018; He et al., 2021; Jiang et al., 2023) incorporated autoregressive structures into entropy models, resulting in more accurate probability predictions. These advancements have laid the foundation for learning-based multi-view image coding.
@@ 3 Proposed Method
Figure 1 shows the overall pipeline of 3D-GP-LMVIC. Given a set of multi-view image sequences $\{\bm{x}_{1},\bm{x}_{2},\bm{x}_{3},\cdots,\bm{x}_{N}\}$, a 3D-GS is trained to estimate depth maps $\bm{d}_{n}$ for each view $\bm{x}_{n}$. Both $\bm{x}_{n}$ and $\bm{d}_{n}$ are compressed, with the coding reference relationships indicated by black solid arrows in Figure 1. Prior to compressing the image $\bm{x}_{n}$, it is necessary to compress $\bm{x}_{n-1}$, $\bm{d}_{n-1}$, and $\bm{d}_{n}$. The disparity relationship between the $(n-1)$-th and $n$-th views is inferred from the reconstructed depth maps $\bm{\hat{d}}_{n-1}$ and $\bm{\hat{d}}_{n}$. Subsequently, based on the estimated disparity relationship, as well as the intermediate features and the reconstructed image $\bm{\hat{x}}_{n-1}$ obtained during the image decoding process of the $(n-1)$-th view, $\bm{x}_{n}$ is compressed. When compressing the depth map $\bm{d}_{n}$, the model employs the predicted depth map derived from $\bm{\hat{d}}_{n-1}$ as a reference. The same neural network architecture and model parameters are used consistently across all views for both image compression and depth map compression. For the initial view lacking reference information, full-zero tensors are provided as the input reference.
The remainder of this section is structured as follows: Section 3.1 elaborates on the method for depth map estimation for a given view using the 3D-GS and the estimation of inter-view disparities. Section 3.2 discusses the compression model for images and depth maps. Section 3.3 introduces a multi-view sequence ordering method designed to ensure sufficient overlap of captured objects between adjacent views.
Figure 2: Illustration of depth-based disparity estimation
@@
For an image $\bm{x}_{n}\in\mathbb{R}^{W\times H\times 3}$ with spatial dimensions $W$ and $H$, we aim to derive a depth map $\bm{d}_{n}\in\mathbb{R}^{W\times H}$, representing the $z$-axis coordinates of each pixel’s corresponding scene points in the camera coordinate system. This depth map facilitates the estimation of disparities between different views.
@@
In Eq. 1, $T_{i}$ serves as a weight for the contribution of each point’s color to the pixel’s final color, diminishing from 1 to 0 as $i$ increases due to cumulative absorption. To estimate the depth of the pixel $d$, we use the depth $z_{i}$ of the first point where $T_{i}$ drops below 0.5, following the median depth estimation approach outlined in Luiten et al. (2024):
@@
Next, we aim to estimate the disparity $\bm{\Delta}_{n}\in\mathbb{R}^{W\times H\times 2}$ between views based on the estimated depth map. This disparity represents the pixel-wise shift of each scene point’s projection across different views. Disparity estimation captures the geometric relationships between views, facilitating the modeling of inter-view correlations.
@@
Figure 2 illustrates the depth-based disparity estimation. To estimate the disparity, a pixel $(x_{n},y_{n})$ in the $n$-th view is back-projected into 3D space using the depth $d_{n}$ to obtain the world coordinates $(x_{\text{w}},y_{\text{w}},z_{\text{w}})$. This 3D world point is then projected into the $(n-1)$-th view to obtain the corresponding pixel coordinates $(x_{n-1},y_{n-1})$. The transformations involved are as follows:
$\begin{bmatrix}x_{\text{w}}\\ y_{\text{w}}\\ z_{\text{w}}\\ 1\end{bmatrix}$ $=V_{n}^{-1}\cdot\text{aug}\left(K^{-1}d_{n}\begin{bmatrix}x_{n}\\ y_{n}\\ 1\end{bmatrix}\right),\hskip 5.0pt\text{with}\hskip 5.0pt\text{aug}\left(\begin{bmatrix}x\\ y\\ z\end{bmatrix}\right)=\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix},$ (3)
@@
$d^{\prime}_{n-1}\begin{bmatrix}x_{n-1}\\ y_{n-1}\\ 1\end{bmatrix}$ $=K\cdot\text{deaug}\left(V_{n-1}\begin{bmatrix}x_{\text{w}}\\ y_{\text{w}}\\ z_{\text{w}}\\ 1\end{bmatrix}\right),\hskip 5.0pt\text{with}\hskip 5.0pt\text{deaug}\left(\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix}\right)=\begin{bmatrix}x\\ y\\ z\end{bmatrix},$
@@
where $K\in\mathbb{R}^{3\times 3}$ denotes the camera intrinsic matrix, and $V_{n},V_{n-1}\in\mathbb{R}^{4\times 4}$ are the extrinsic matrices corresponding to the $n$-th and $(n-1)$-th views, respectively. The camera parameters are calibrated using SfM (Schonberger & Frahm, 2016). $d^{\prime}_{n-1}$ represents the depth of the 3D world point in the camera coordinate system of the $(n-1)$-th view. The resulting disparity $\delta_{n}=(x_{n-1}-x_{n},y_{n-1}-y_{n})$ for each pixel is then compiled into the disparity map $\bm{\Delta}_{n}$.
Finally, we define a mask $\bm{x}_{n,\text{m}}\in\mathbb{R}^{W\times H}$ that identifies whether each pixel in the $n$-th view maps to the same 3D world point as the corresponding pixel in the $(n-1)$-th view through disparity estimation. The mask’s criteria are: (1) the projected pixel must reside within the valid image region in the $(n-1)$-th view, (2) the corresponding 3D world point must be in the positive $z$-half-space of the $(n-1)$-th view’s coordinate system, and (3) no occlusion must exist along the line of sight, i.e., $d^{\prime}_{n-1}$ from Eq. 3 must be less than the estimated depth along the ray in the $(n-1)$-th view. This can be formulated as:
@@
$\bm{x}_{n,\text{m}}[i,j]=\begin{cases}1&\text{if}\hskip 5.0pt0<\bm{\Delta}_{n}[i,j,0]+i+0.5<W\\ &\hskip 5.0pt\text{and}\hskip 5.0pt0<\bm{\Delta}_{n}[i,j,1]+j+0.5<H\\ &\hskip 5.0pt\text{and}\hskip 5.0pt0<\bm{d}^{\prime}_{n-1}[i,j]<\text{Warp}(\bm{d}_{n-1},\bm{\Delta}_{n})[i,j],\\ 0&\text{otherwise},\end{cases}$ (4)
@@
where $\bm{d}^{\prime}_{n-1}\in\mathbb{R}^{W\times H}$ represents the tensor containing the depth values $d^{\prime}_{n-1}$ for each pixel, and $\text{Warp}(\cdot,\cdot)$ denotes the warping operation based on the given disparity.
Figure 3: The proposed image compression model. Convolution and deconvolution layers are specified by the tuple (number of output channels, kernel size, stride). ’Q’ denotes the quantization step. ’AE’ and ’AD’ stand for arithmetic encoder and arithmetic decoder, respectively.

### v2

# 3D-LMVIC: Learning-based Multi-View Image Compression with 3D Gaussian Geometric Priors

@@ Abstract
Existing multi-view image compression methods often rely on 2D projection-based similarities between views to estimate disparities. While effective for small disparities, such as those in stereo images, these methods struggle with the more complex disparities encountered in wide-baseline multi-camera systems, commonly found in virtual reality and autonomous driving applications. To address this limitation, we propose 3D-LMVIC, a novel learning-based multi-view image compression framework that leverages 3D Gaussian Splatting to derive geometric priors for accurate disparity estimation. Furthermore, we introduce a depth map compression model to minimize geometric redundancy across views, along with a multi-view sequence ordering strategy based on a defined distance measure between views to enhance correlations between adjacent views. Experimental results demonstrate that 3D-LMVIC achieves superior performance compared to both traditional and learning-based methods. Additionally, it significantly improves disparity estimation accuracy over existing two-view approaches.
Machine Learning, ICML
@@ 1 Introductioin
Figure 1: Illustrations of camera systems. (a) A stereo camera configuration. (b) A wide-baseline multi-camera configuration.
@@
The rapid advancement of 3D applications has led to an explosion of multi-view image data across various fields, including virtual reality (VR) (Anthes et al., 2016), augmented reality (AR) (Schmalstieg & Hollerer, 2016), visual simultaneous localization and mapping (vSLAM) (Mokssit et al., 2023), 3D scene understanding (Dai et al., 2017), autonomous driving (Chen et al., 2017), and medical imaging (Hosseinian & Arefi, 2015). In particular, applications like VR and AR, which rely on high-quality multi-view visual content to create immersive experiences, generate a massive volume of data that poses significant challenges for storage and transmission. This makes the development of efficient compression techniques crucial for managing the increasing data demands in these fields.
As illustrated in Figure 1, unlike the commonly studied stereo camera systems, 3D applications often rely on wide-baseline multi-camera systems to capture global scene information (Xu et al., 2020; Yan et al., 2024). In such scenarios, the spatial positions and viewing angles of cameras differ significantly compared to stereo setups, resulting in large disparities between images captured from different views. Existing disparity estimation methods typically rely on finding similar local regions in the image domain to estimate disparities. However, this approach faces significant challenges when dealing with complex and large disparities.
@@
Current multi-view coding standards, such as H.264-based MVC (Vetro et al., 2011) and H.265-based MV-HEVC (Hannuksela et al., 2015), have been developed to compress multi-view media by extending their respective base standards and exploiting redundancies across multiple views. These standards employ disparity estimation to calculate positional differences of objects between views, aiding in the prediction of pixel values. However, these methods rely on manually designed modules, limiting the system’s ability to fully leverage end-to-end optimization.
@@
Learning-based single image compression has seen remarkable advancements (Ballé et al., 2017, 2018; Minnen et al., 2018), inspiring extensions of these methods to multi-view image coding (Deng et al., 2021; Lei et al., 2022; Zhang et al., 2023; Liu et al., 2024). A central challenge in these extensions lies in the accurate estimation of disparities across different views. For example, Deng et al. (2021, 2023) employ a simple 3x3 homography matrix for disparity estimation, which, while efficient, struggles with complex scene disparities. Alternatively, Ayzik & Avidan (2020); Huang et al. (2023) utilize patch matching method to align the reference view with the target view. This approach is effective for horizontal or vertical view shifts but falls short when addressing non-rigid deformations caused by view rotations. Similarly, Zhai et al. (2022) assume that disparity occurs only along the horizontal axis in their stereo matching method, which suffices for stereo images but is inadequate for more complex view transformations where disparity is not limited to the horizontal axis. Some methods leverage cross-attention mechanisms for implicit alignment (Wödlinger et al., 2022; Zhang et al., 2023; Liu et al., 2024). For instance, Zhang et al. (2023) enhance the target view’s representation by multiplying its query with the reference view’s key and value, effectively incorporating reference view features into the target view. However, these methods primarily establish correlations between two views by 2D projection similarities, without considering the 3D spatial relationships between the views and the captured objects.
@@
Building on prior investigation, we propose a novel learning-based multi-view image compression framework with 3D Gaussian geometric priors (3D-LMVIC), which employs 3D-GS as a geometric prior to guide disparity estimation between views. Specifically, 3D-GS generates a depth map for each view, providing precise spatial information at the pixel level. This enables accurate correspondence between views, allowing the compression model to effectively fuse features from reference views. Due to positional and angular disparities between views, images generally do not fully overlap, and merging non-overlapping regions may introduce noise. To address this, we design a mask based on the 3D Gaussian geometric prior to identify overlapping regions, ensuring more accurate feature fusion. Additionally, since depth maps are required during decoding, we propose a depth map compression model to efficiently reduce geometric redundancy across views, incorporating a cross-view depth prediction module to capture inter-view geometric correlations. Finally, recognizing the importance of field of view (FoV) overlap in redundancy reduction, we introduce a multi-view sequence ordering method to address the issue of low overlap between adjacent views in unordered sequences. This method defines and proves a distance measure between view pairs to guide the ordering of view sequences.
@@
  • We propose a learning-based multi-view image compression framework with 3D Gaussian geometric priors (3D-LMVIC), which utilizes 3D Gaussian geometric priors for precise disparity estimation between views, thereby enhancing multi-view image compression efficiency. Additionally, we design a mask based on these priors to identify overlapping regions between views, effectively guiding the model to retain useful cross-view information.
@@
  • We also present a depth map compression model aimed at reducing geometric redundancy across views. Additionally, we define and prove a distance measure between views, upon which a multi-view sequence ordering method is proposed to improve the correlation between adjacent views.
@@
  • Experimental results show that our framework surpasses both traditional and learning-based multi-view image coding methods in compression efficiency. Moreover, our disparity estimation method demonstrates greater visual accuracy compared to existing two-view disparity estimation methods.
@@ 2 Related Works
In recent years, learning-based image compression methods have integrated autoencoders with differentiable entropy models to enable end-to-end optimization of rate-distortion loss. Early works, such as Ballé et al. (2017, 2018), introduced generalized divisive normalization (GDN) (Ballé et al., 2016) and proposed factorized and hyperprior entropy models. Subsequent research (Minnen et al., 2018; He et al., 2021; Jiang et al., 2023) incorporated autoregressive structures into entropy models, resulting in more accurate probability predictions. These advancements have laid the foundation for learning-based multi-view image coding.
@@
(a) Overview of the 3D-LMVIC pipeline.
(b) Depth-based disparity estimation process.
Figure 2: Overall Pipeline.
@@ 3 Proposed Method
Figure 2(a) shows the overall pipeline of 3D-LMVIC. Given a set of multi-view image sequences $\mathcal{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\boldsymbol{x}_{3},\cdots,\boldsymbol{x}_{N}\}$, a 3D-GS is trained to estimate depth map $\boldsymbol{d}_{n}$ for each image $\boldsymbol{x}_{n}$. Both $\boldsymbol{x}_{n}$ and $\boldsymbol{d}_{n}$ are compressed, with the coding reference relationships indicated by black solid arrows in the figure. Prior to compressing the image $\boldsymbol{x}_{n}$, it is necessary to compress $\boldsymbol{x}_{n-1}$, $\boldsymbol{d}_{n-1}$, and $\boldsymbol{d}_{n}$. The disparity relationship between the $(n-1)$-th and $n$-th views is inferred from the reconstructed depth maps $\boldsymbol{\hat{d}}_{n-1}$ and $\boldsymbol{\hat{d}}_{n}$. Subsequently, based on the estimated disparity relationship, as well as the extracted features of the $(n-1)$-th view, $\boldsymbol{x}_{n}$ is compressed. When compressing the depth map $\boldsymbol{d}_{n}$, the model employs the predicted depth map derived from $\boldsymbol{\hat{d}}_{n-1}$ as a reference. The same neural network architecture and model parameters are used consistently across all views for both image compression and depth map compression.
The remainder of this section is structured as follows: Section 3.1 elaborates on the method for depth map estimation for a given view using the 3D-GS and the estimation of inter-view disparities. Section 3.2 covers the compression model for both images and depth maps, as well as the multi-view sequence ordering method.
@@
3.1.1 Depth Estimation
For an image $\boldsymbol{x}_{n}\in\mathbb{R}^{W\times H\times 3}$ with spatial dimensions $W$ and $H$, we aim to derive a depth map $\boldsymbol{d}_{n}\in\mathbb{R}^{W\times H}$, representing the $z$-axis coordinates of each pixel’s corresponding 3D world point in the camera coordinate system. This depth map facilitates the estimation of disparities between different views.
@@ 3.1.1 Depth Estimation
In (1), $T_{i}$ serves as a weight for the contribution of each point’s color to the pixel’s final color, diminishing from 1 to 0 as $i$ increases due to cumulative absorption. To estimate the depth of a pixel $d$, we adopt a median depth estimation approach. Specifically, the depth is determined as the depth of the first point where $T_{i}$ drops below 0.5:
@@
It is worth noting that the original 3D-GS (Kerbl et al., 2023) employs a weighted averaging approach, using $T_{i}\alpha_{i}$ as the weight for each 3D Gaussian along the ray to compute depth. In contrast, alignment experiments in Section 4.3 demonstrate that the median depth estimation approach achieves better alignment performance.
3.1.2 Disparity Estimation
@@
Next, we aim to estimate the disparity $\boldsymbol{\Delta}_{n}\in\mathbb{R}^{W\times H\times 2}$ between views based on the estimated depth map. This disparity represents the pixel-wise shift of each 3D world point’s projection across different views. Disparity estimation captures the geometric relationships between views, facilitating the modeling of inter-view correlations.
@@
Figure 2(b) illustrates the depth-based disparity estimation. To estimate the disparity, a pixel $(x_{n},y_{n})$ in the $n$-th view is back-projected into 3D space using the depth $d_{n}$ to obtain the world coordinates $(x_{\text{w}},y_{\text{w}},z_{\text{w}})$. This 3D world point is then projected into the $(n-1)$-th view to obtain the corresponding pixel coordinates $(x_{n-1},y_{n-1})$. The transformations involved are as follows:
@@
$\begin{bmatrix}x_{\text{w}}\\ y_{\text{w}}\\ z_{\text{w}}\\ 1\end{bmatrix}&=V_{n}^{-1}\cdot\text{aug}\left(K^{-1}d_{n}\begin{bmatrix}x_{n}\\ y_{n}\\ 1\end{bmatrix}\right),\\ d^{\prime}_{n-1}\begin{bmatrix}x_{n-1}\\ y_{n-1}\\ 1\end{bmatrix}&=K\cdot\text{deaug}\left(V_{n-1}\begin{bmatrix}x_{\text{w}}\\ y_{\text{w}}\\ z_{\text{w}}\\ 1\end{bmatrix}\right),\end{aligned}$ (3)
@@
where $K\in\mathbb{R}^{3\times 3}$ denotes the camera intrinsic matrix, and $V_{n},V_{n-1}\in\mathbb{R}^{4\times 4}$ are the extrinsic matrices corresponding to the $n$-th and $(n-1)$-th views, respectively. The camera parameters are calibrated using SfM (Schonberger & Frahm, 2016). $d^{\prime}_{n-1}$ represents the depth of the 3D world point in the camera coordinate system of the $(n-1)$-th view. aug denotes the operation of augmenting a vector by adding an additional dimension with a value of 1 as its final element. Conversely, deaug refers to the operation of reducing a vector by removing its last dimension. The resulting disparity $\delta_{n}=(x_{n-1}-x_{n},y_{n-1}-y_{n})$ for each pixel is then compiled into the disparity map $\boldsymbol{\Delta}_{n}$.
Finally, we define a mask $\boldsymbol{x}_{n,\text{m}}\in\mathbb{R}^{W\times H}$ to determine whether the disparity estimation is meaningful, i.e., whether the information from the reference pixel is relevant or merely noise. The mask’s criteria are as follows:
1. The projected pixel must reside within the valid image region in the $(n-1)$-th view.
2. The corresponding 3D world point must lie in the positive $z$-half-space of the $(n-1)$-th view’s coordinate system.
3. No occlusion must exist along the line of sight, i.e., $d^{\prime}_{n-1}$ from (3) must be less than the estimated depth along the ray in the $(n-1)$-th view.
This can be formulated as:
@@
$\boldsymbol{x}_{n,\text{m}}[i,j]=\begin{cases}1&\text{if}\hskip 5.0pt0<\boldsymbol{\Delta}_{n}[i,j,0]+i+0.5<W\\ &\text{and}\hskip 5.0pt0<\boldsymbol{\Delta}_{n}[i,j,1]+j+0.5<H\\ &\text{and}\hskip 5.0pt0<\boldsymbol{d}^{\prime}_{n-1}[i,j]<\text{Warp}(\boldsymbol{d}_{n-1},\boldsymbol{\Delta}_{n})[i,j],\\ 0&\text{otherwise},\end{cases}$ (4)
@@
where $\boldsymbol{d}^{\prime}_{n-1}\in\mathbb{R}^{W\times H}$ represents the tensor containing the depth values $d^{\prime}_{n-1}$ for each pixel, and $\text{Warp}(\cdot,\cdot)$ denotes the warping operation based on the given disparity. Appendix A outlines the algorithmic process for disparity and mask estimation.
Figure 3: The architecture of the proposed image compression model. ’LR’ represents the Leaky ReLU activation function, ’Q’ denotes the quantization operation, and ’AE’/’AD’ refer to the arithmetic encoder/decoder, respectively.